# Blessing of Multilinguality: A Systematic Analysis of Multilingual In-Context Learning

### v1

@@ 4.3 Ablation Study: Merely Introducing New Language(s) Enhances ICL Performance
After showing that including non-English in the prompts can improve ICL performance, a natural follow-up question arises: does the gain come from the mere presence of non-English languages, or the interaction between the target language and the in-topic examples? To distinguish these two setups, we prepend a context-irrelevant sentence (CIS) before each ICL demonstration, which is unrelated to the current domain and can be in any language. Analogous to the ICL modes introduced in § 2.2, CIS resembles the construction of those modes with the same sampling strategy (§ 3). For example, based on the English mode, we could prepend a set of multilingual CIS, which augments Fig. 2a into Fig. 5. We denote such setting by “English CIS-Multi”. This naming convention applies to other settings accordingly.
@@
We use sentences from FLORES-101 Goyal et al. (2022) as the source of our CIS, which provides parallel Wikipedia sentences in multiple languages, a fairly distant genre to all evaluation datasets. We filter sentence-parallel sets with English word counts ranging from 10 to 15 as our sampling pool , a small range compared to target datasets (Tab. 1), to mitigate the risk of introducing too much noise. Details of FLORES-101 can be found at § A.2.
@@
Results and analysis. We perform hypothesis testing as in Tab. 2 to compare different CIS settings with English CIS-En (Tab. 3)—English CIS-En exhibit a small drop compared to English only, indicating that our filtration effectively controls the negative impact of noise to a tolerable level. We find that introducing a single non-English language generally improves ICL performance in most cases ; however, the improvement is only statistically significant in cases. This reveals that simply introducing a language can lead to a modest improvement in MLLM performance, but it is more pronounced when in-topic demonstrations in another language (Monolingual modes) are incorporated. More hypothesis test results for both LRL and HRL splits can be found in Tabs. 20, 21 and 22 in § B.2.
@@
Introducing multiple languages (CIS-Multi, Fig. 5) is slightly more promising than CIS in a single language, with of cases showing improvement, of which are significant. We further conduct experiments by prepending multilingual CIS to the Multilingual mode (Multilingual CIS-Multi, Tab. 3). The performance of Multilingual CIS-Multi is not significantly lower than that of English CIS-Multi; in more than half cases, it significantly outperforms English CIS-Multi. This concludes that the significant improvement from English mode to Multilingual mode is attributed to the incorporation of multiple languages with in-topic demonstrations.
| LRL AvgΔ↑↓ | MGSM | XL-WiC | XCOPA |
| English | 64.20 | 57.37 | 46.23 |
| English CIS-En | 61.00 | 56.73 | 48.37 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
| English | 57.10 | 44.42 | 55.91 |
| English CIS-En | 55.90 | 47.88 | 55.46 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
| English | 43.70 | 48.46 | 62.29 |
| English CIS-En | 43.00 | 50.90 | 62.31 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
| English | 59.40 | 57.82 | 64.69 |
| English CIS-En | 59.40 | 59.04 | 64.20 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
| English | 57.00 | 51.54 | 62.26 |
| English CIS-En | 60.70 | 49.62 | 61.00 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
| English | 29.40 | 58.78 | 26.54 |
| English CIS-En | 27.80 | 57.82 | 23.20 |
| English CIS-Fr | |||
| English CIS-Ja | |||
| English CIS-Zh | |||
| English CIS-Multi | |||
| Multilingual CIS-Multi | |||
@@ 4.4 Translation-Based Performance
Strategies. We test two translation strategies for baseline comparison. (1) Transl-LangEn: Translating test questions in other languages in the English ICL mode (Fig. 2a) into English. (2) Transl-EnLang: Translating demonstrations in English ICL mode (Fig. 2a) into the source language of the current test question, which mirrors the Native mode (Fig. 2d). We use the Google Cloud Translation API for translation.333https://cloud.google.com/translate/
@@
Analysis. For LRLs, Transl-LangEn sometimes outperform the Native mode, while Transl-EnLang performs comparably to the Multilingual mode, but falls short of the Native mode performance (Tab. 4). These results resonate with the phenomena that the translation quality of LRLEn is generally higher than that of the reversed direction Fan et al. (2021); Goyal et al. (2022); Costa-jussà et al. (2022).
@@ 5 Related Work
Prompt engineering. Instruction tuning aligns LLMs more closely with human instructions Ouyang et al. (2022); Mishra et al. (2022); Wei et al. (2022a); Askell et al. (2021); Wang et al. (2022, 2023). Concurrently, numerous prompting strategies have been developed Liu et al. (2023) and shown to consistently enhance the performance of instruction-tuned LLMs, such as in-context learning Brown et al. (2020); Min et al. (2022) and chain-of-thought Wei et al. (2022b); Kojima et al. (2022). These prompting strategies are proven effective in multilingual tasks as well Winata et al. (2021); Lin et al. (2022); Shi et al. (2023).
Multilingual ICL. For languages of templates, demonstration and sample question in native language is conventionally inserted into a predefined English template Lin et al. (2022); Fu et al. (2022); Qin et al. (2023); Huang et al. (2023); Ahuja et al. (2023); Zhang et al. (2024b). Poelman and de Lhoneux (2024) critiques this widespread misuse of English as the interface language. Qin et al. (2023); Huang et al. (2023); Zhang et al. (2024b) guide models to “think” and generate CoT in English, regardless of the input language, leading to improved performance for generation tasks compared to “thinking” in other language(s). Sclar et al. (2024); Zhang et al. (2024a) highlight that models are sensitive to those templates. For languages of demonstrations and test questions, Shi et al. (2023); Ahuja et al. (2023) conclude that in-language demonstrations outperform English demonstrations. Etxaniz et al. (2024); Liu et al. (2024) suggest translating questions from LRLs into English can improve performance.
@@ 6 Conclusion and Discussion
This paper treats multilingual LLMs as black-box models, drawing the findings and conclusions based solely on their input-output behavior. Hence we have not interpreted the internal mechanism how multilinguality could affect MLLM’s “thinking” process and its manifested performance. We have briefly touched the impact of demonstrations in different languages on the MLLM performance. However, we do not conduct a thorough empirical analysis to identify which specific linguistic characteristics (e.g., writing systems, grammatical structures, or linguistic relatedness) contribute to the observed performance differences.
Reproducibility
Our code is available at https://github.com/yileitu/multilingual_icl.
@@ A.1 Model
[…]
All model checkpoints we use and their properties are listed in Tab. 5.
@@ A.2 Dataset
After preprocessing, datapoints for each language split are stored in a single JSON file. Tab. 6 summarizes the supported languages for each dataset and the sources from which they are obtained.
@@
The original dataset consists of parallel datapoints across all language splits and training/test splits of the same size. Example datapoint and the Chat Template for few-shot demonstrations can be found in Fig. 6.
@@
The datapoints in the training split of XCOPA are parallel to the last 100 datapoints in the English COPA development split. Therefore, we exclude the first datapoints from the development split of English COPA. The test splits of both datasets are parallel and contain the same number of datapoints. We then merge both into our XCOPA dataset. In XCOPA, there are two types of questions: “cause” and “effect”, each corresponding to a distinct template, as shown in Figs. 7(a) and 7(b).
@@
This benchmark is designed to determine whether a specific word in a given language has the same meaning in two different sentences. As a result, the dataset is inherently non-parallel. Among all language splits, Estonian (et) contains the fewest datapoints, with in the training split and in the test split. For all other languages, we randomly subsample to match the size of the Estonian split to satisfy the demonstration sampling requirements outlined in Fig. 2. To leverage the attention mechanism of transformers Vaswani et al. (2017), we add asterisks around the target word in both sentences to indicate that the LLM needs to disambiguate the meaning of that specific word. An example datapoint is shown in Fig. 8.
@@
This machine translation benchmark comprises sentences extracted from English Wikipedia, spanning diverse topics and domains. These sentences were translated into languages by professional translators via a carefully controlled process. The Wikipedia domain is largely unrelated to the domains of the three datasets we evaluate (math, commonsense reasoning, and word disambiguation). Therefore, we choose FLORES-101 as our source pool of irrelevant sentences. Note that in this benchmark, each datapoint carries the same semantic meaning across all language splits. We do not want irrelevant sentences to affect the LLM’s understanding of the original task excessively, thus we select sentences with word counts between and in the English split, introducing limited noise. CIS are drawn form the filtered FLORES-101 dataset. An example is provided in Fig. 9.
@@ A.3.1 High-Resource Language List
PaLM Chowdhery et al. (2023)—that publicly report the language distribution used during pretraining. The top 20 languages and their percentages for each model are listed in Tab. 7 and Tab. 8, respectively. We take the union of these languages to form what we consider a high-resource language list (in terms of the richness in the LLM pretraining dataset), including the following languages:
@@ A.3.2 Languages We Evaluate
Tab. 9 presents the union of languages supported by the three datasets we evaluated (§ 3, datasets). Based on whether a language appears in the high-resource language list, we categorized the languages into HRL and LRL groups, with classified as HRL and as LRL. Among them, only English and Chinese are present in all three datasets.
@@
| bg | Bulgarian | Low | XL-WiC | Cyrillic | Indo-European |
| bn | Bengali | Low | MGSM | Bengali–Assamese | Indo-European |
| da | Danish | High | XL-WiC | Latin | Indo-European |
| de | German | High | MGSM, XL-WiC | Latin | Indo-European |
| en | English | High | MGSM, XL-WiC, XCOPA | Latin | Indo-European |
| es | Spanish | High | MGSM | Latin | Indo-European |
| et | Estonian | Low | XCOPA | Latin | Indo-European |
| fa | Persian | Low | XL-WiC | Arabic | Indo-European |
| fr | French | High | MGSM, XL-WiC | Latin | Indo-European |
| hr | Croatian | Low | XL-WiC | Latin | Indo-European |
| ht | Haitian | Low | XCOPA | Latin | French Creole |
| id | Indonesian | High | XCOPA | Latin | Austronesian |
| it | Italian | High | XL-WiC, XCOPA | Latin | Indo-European |
| ja | Japanese | High | MGSM, XL-WiC | Kana & Chinese Characters | Japonic |
| ko | Korean | High | XL-WiC | Hangul | Koreanic |
| nl | Dutch | High | XL-WiC | Latin | Indo-European |
| qu | Quechua | Low | XCOPA | Latin | Quechumaran |
| ru | Russian | High | MGSM | Cyrillic | Indo-European |
| sw | Swahili | Low | MGSM | Latin | Niger–Congo |
| ta | Tamil | Low | XCOPA | Tamil | Dravidian |
| te | Telegu | Low | MGSM | Telegu | Dravidian |
| th | Thai | Low | MGSM, XCOPA | Thai | Kra–Dai |
| tr | Turkish | High | XCOPA | Latin | Turkic |
| vi | Vietnamese | Low | XCOPA | Latin | Austroasiatic |
| zh | Chinese | High | MGSM, XL-WiC, XCOPA | Chinese Characters | Sino-Tibetan |
@@ C.1.1 Identifying Top-Activated Neurons
Each neuron in every MLP layer of the model is assigned a dedicated counter, initialized to . During vanilla evaluation, we monitor the activation of every neuron in the forward pass. Since LLMs typically employ ReLU-like Agarap (2018) activation functions (e.g., SwiGLU Shazeer (2020) for Llama series), a positive activation value can be interpreted as the neuron being “activated”. If a neuron is “activated”, we increment the corresponding counter by , otherwise no action. To ensure balanced input across our three datasets, we curated a COMBINED dataset, see § A.2 for details. After processing the inputs of a single ICL mode, each neuron accumulates an “activated” count. The neurons with the highest counts are identified as specialized neurons attributed to this ICL mode.

### v3

@@ 4.3 Ablation Study: Merely Introducing New Language(s) Enhances ICL Performance
After showing that including non-English in the prompts can improve ICL performance, a natural follow-up question arises: does the gain come from the mere presence of non-English languages, or the interaction between the target language and the in-topic examples? To distinguish these two setups, we prepend a context-irrelevant sentence (CIS) before each ICL demonstration, which is unrelated to the current domain and can be in any language. Analogous to the ICL modes introduced in §˜2.2, CIS resembles the construction of those modes with the same sampling strategy (§˜3). For example, based on the English mode, we could prepend a set of multilingual CIS, which augments Fig.˜2a into Fig.˜5. We denote such setting by “English CIS-Multi”. This naming convention applies to other settings accordingly.
@@
We use sentences from FLORES-101 Goyal et al. (2022) as the source of our CIS, which provides parallel Wikipedia sentences in multiple languages, a fairly distant genre to all evaluation datasets. We filter sentence-parallel sets with English word counts ranging from 10 to 15 as our sampling pool , a small range compared to target datasets (Tab.˜1), to mitigate the risk of introducing too much noise. Details of FLORES-101 can be found at §˜A.2.
@@
Results and analysis. We perform hypothesis testing as in Tab.˜2 to compare different CIS settings with English CIS-En (Tab.˜3)—English CIS-En exhibit a small drop compared to English only, indicating that our filtration effectively controls the negative impact of noise to a tolerable level. We find that introducing a single non-English language generally improves ICL performance in most cases ; however, the improvement is only statistically significant in cases. This reveals that simply introducing a language can lead to a modest improvement in MLLM performance, but it is more pronounced when in-topic demonstrations in another language (Monolingual modes) are incorporated. More hypothesis test results for both LRL and HRL splits can be found in Tabs.˜23, 24 and 25 in §˜B.2.
@@
Introducing multiple languages (CIS-Multi, Fig.˜5) is slightly more promising than CIS in a single language, with of cases showing improvement, of which are significant. We further conduct experiments by prepending multilingual CIS to the Multilingual mode (Multilingual CIS-Multi, Tab.˜3). The performance of Multilingual CIS-Multi is not significantly lower than that of English CIS-Multi; in cases, it significantly outperforms English CIS-Multi. This concludes that the significant improvement from English mode to Multilingual mode is attributed to the incorporation of multiple languages with in-topic demonstrations.
| LRL | MGSM | XL-WiC | XCOPA | XQuAD |
| English | 64.20 | 57.37 | 46.23 | 72.02 |
| CIS-En | 61.00 | 56.73 | 48.37 | 71.43 |
| CIS-Fr | ||||
| CIS-Ja | ||||
| CIS-Zh | ||||
CIS- |
| English | 57.10 | 44.42 | 55.91 | 68.25 |
| CIS-En | 55.90 | 47.88 | 55.46 | 68.45 |
| CIS-Fr | ||||
| CIS-Ja | ||||
| CIS-Zh | ||||
CIS- |
| English | 43.70 | 48.46 | 62.29 | 53.40 |
| CIS-En | 43.00 | 50.90 | 62.31 | 58.55 |
| CIS-Fr | ||||
| CIS-Ja | ||||
| CIS-Zh | ||||
CIS- |
| English | 59.40 | 57.82 | 64.69 | 68.98 |
| CIS-En | 59.40 | 59.04 | 64.20 | 69.05 |
| CIS-Fr | ||||
| CIS-Ja | ||||
| CIS-Zh | ||||
CIS- |
| English | 57.00 | 51.54 | 62.26 | 60.93 |
| CIS-En | 60.70 | 49.62 | 61.00 | 56.48 |
| CIS-Fr | ||||
| CIS-Ja | ||||
| CIS-Zh | ||||
CIS- |
| English | 29.40 | 58.78 | 26.54 | 70.48 |
| CIS-En | 27.80 | 57.82 | 23.20 | 70.30 |
| CIS-Fr | ||||
| CIS-Ja | ||||
| CIS-Zh | ||||
CIS- |
@@ 4.4 Translation-Based Performance
Strategies. We test two translation strategies for baseline comparison. (1) Transl-LangEn: Translating test questions in other languages in the English ICL mode (Fig.˜2a) into English. (2) Transl-EnLang: Translating demonstrations in English ICL mode (Fig.˜2a) into the source language of the current test question, which mirrors the Native mode (Fig.˜2d). We use the Google Cloud Translation API for translation.222https://cloud.google.com/translate/
@@
Analysis. For LRLs, Transl-LangEn sometimes outperform the Native mode, while Transl-EnLang performs comparably to the Multilingual mode, but falls short of the Native mode performance (Tab.˜4). These results resonate with the phenomena that the translation quality of LRLEn is generally higher than that of the reversed direction Fan et al. (2021); Goyal et al. (2022); Costa-jussà et al. (2022).
@@ 5 Related Work
Prompt engineering. Instruction tuning aligns LLMs more closely with human instructions Ouyang et al. (2022); Mishra et al. (2022); Wei et al. (2022a); Askell et al. (2021); Wang et al. (2022, 2023). Concurrently, numerous prompting strategies have been developed Liu et al. (2023) and shown to consistently enhance the performance of instruction-tuned LLMs, such as in-context learning Brown et al. (2020); Min et al. (2022) and chain-of-thought Wei et al. (2022b); Kojima et al. (2022). These prompting strategies are proven effective in multilingual tasks as well Winata et al. (2021); Lin et al. (2022); Shi et al. (2023). Concurrent work Gao et al. (2025) shows that multilingual prompting has improved the general reasoning capabilities of multilingual LLMs.
Multilingual ICL. For languages of templates, demonstrations and sample questions in native languages are conventionally inserted into a predefined English template Lin et al. (2022); Fu et al. (2022); Qin et al. (2023); Huang et al. (2023); Ahuja et al. (2023); Zhang et al. (2024b). Poelman and de Lhoneux (2024) critiques this widespread misuse of English as the interface language. Qin et al. (2023); Huang et al. (2023); Zhang et al. (2024b) guide models to “think” and generate CoT in English, regardless of the input language, leading to improved performance for generation tasks compared to “thinking” in other language(s). Sclar et al. (2024); Zhang et al. (2024a) highlight that models are sensitive to those templates. For languages of demonstrations and test questions, Shi et al. (2023); Ahuja et al. (2023) conclude that in-language demonstrations outperform English demonstrations. Etxaniz et al. (2024); Liu et al. (2024) suggest translating questions from LRLs into English can improve performance.
@@ 6 Conclusion and Discussion
This paper treats multilingual LLMs as black-box models, drawing the findings and conclusions based solely on their input-output behavior. Hence, we have not interpreted the internal mechanism of how multilingualism could affect MLLM’s “thinking” process and its manifested performance. We have briefly touched on the impact of demonstrations in different languages on the MLLM performance. However, we do not conduct a thorough empirical analysis to identify which specific linguistic characteristics (e.g., writing systems, grammatical structures, or linguistic relatedness) contribute to the observed performance differences.
@@ A.1 Model
8B | meta-llama/Meta-Llama-3-8B-Instruct | ||
8B | meta-llama/Meta-Llama-3.1-8B-Instruct | ||
7B | Qwen/Qwen2-7B-Instruct | ||
7B | Qwen/Qwen2.5-7B-Instruct | ||
12B | mistralai/Mistral-Nemo-Instruct-2407 | ||
8B | CohereForAI/aya-expanse-8b | ||
NA | gpt-3.5-turbo-0125 | ||
NA | gpt-4o-mini-2024-07-18 |
All model checkpoints we use and their properties are listed in Tab.˜5.
@@ A.2 Dataset
| XQuAD | ar, de, en, es, ru, tr, vi, zh | el, hi, ro, th | google/xquad |
After preprocessing, datapoints for each language split are stored in a single JSON file. Tab.˜6 summarizes the supported languages for each dataset and the sources from which they are obtained.
@@
The original dataset consists of parallel datapoints across all language splits and training/test splits of the same size. Example datapoint and the Chat Template for few-shot demonstrations can be found in Fig.˜6.
@@
The datapoints in the training split of XCOPA are parallel to the last 100 datapoints in the English COPA development split. Therefore, we exclude the first datapoints from the development split of English COPA. The test splits of both datasets are parallel and contain the same number of datapoints. We then merge both into our XCOPA dataset. In XCOPA, there are two types of questions: “cause” and “effect”, each corresponding to a distinct template, as shown in Figs.˜7(a) and 7(b).
@@
This benchmark is designed to determine whether a specific word in a given language has the same meaning in two different sentences. As a result, the dataset is inherently non-parallel. Among all language splits, Estonian (et) contains the fewest datapoints, with in the training split and in the test split. For all other languages, we randomly subsample to match the size of the Estonian split to satisfy the demonstration sampling requirements outlined in Fig.˜2. To leverage the attention mechanism of transformers Vaswani et al. (2017), we add asterisks around the target word in both sentences to indicate that the LLM needs to disambiguate the meaning of that specific word. An example datapoint is shown in Fig.˜8.
@@
XQuAD
The dataset consists of parallel question-answering examples across languages, derived from the same subset of the English SQuAD Rajpurkar et al. (2016) v1.1 development set. For each language, the context passages, questions, and answers are professionally translated while preserving semantic alignment across languages. Each example includes a context paragraph, a question, and a span-based answer. An example and the input prompt format used in our experiments are illustrated in Fig.˜9.
@@
This machine translation benchmark comprises sentences extracted from English Wikipedia, spanning diverse topics and domains. These sentences were translated into languages by professional translators via a carefully controlled process. The Wikipedia domain is largely unrelated to the domains of the three datasets we evaluate (math, commonsense reasoning, and word disambiguation). Therefore, we choose FLORES-101 as our source pool of irrelevant sentences. Note that in this benchmark, each datapoint carries the same semantic meaning across all language splits. We do not want irrelevant sentences to affect the LLM’s understanding of the original task excessively, thus we select sentences with word counts between and in the English split, introducing limited noise. CIS are drawn form the filtered FLORES-101 dataset. An example is provided in Fig.˜10.
@@ A.3.1 High-Resource Language List
PaLM Chowdhery et al. (2023)—that publicly report the language distribution used during pretraining. The top 20 languages and their percentages for each model are listed in Tab.˜7 and Tab.˜8, respectively. We take the union of these languages to form what we consider a high-resource language list (in terms of the richness in the LLM pretraining dataset), including the following languages:
@@ A.3.2 Languages We Evaluate
Tab.˜9 presents the union of languages supported by the datasets we evaluated (Tab.˜1). Based on whether a language appears in the high-resource language list, we categorized the languages into HRL and LRL groups, with classified as HRL and as LRL. Among them, only English and Chinese are present in all three datasets.
@@
| ar | Arabic | H | XQuAD | Arabic | Afro-Asiatic |
| bg | Bulgarian | L | XL-WiC | Cyrillic | Indo-European |
| bn | Bengali | L | MGSM | Bengali–Assamese | Indo-European |
| da | Danish | H | XL-WiC | Latin | Indo-European |
| de | German | H | MGSM, XL-WiC, XQuAD | Latin | Indo-European |
| el | Greek | L | XQuAD | Greek | Indo-European |
| en | English | H | MGSM, XCOPA, XL-WiC XQuAD | Latin | Indo-European |
| es | Spanish | H | MGSM, XQuAD | Latin | Indo-European |
| et | Estonian | L | XCOPA | Latin | Indo-European |
| fa | Persian | L | XL-WiC | Arabic | Indo-European |
| fr | French | H | MGSM, XL-WiC | Latin | Indo-European |
| hi | Hindi | L | XQuAD | Devanagari | Indo-European |
| hr | Croatian | L | XL-WiC | Latin | Indo-European |
| ht | Haitian | L | XCOPA | Latin | French Creole |
| id | Indonesian | H | XCOPA | Latin | Austronesian |
| it | Italian | H | XL-WiC, XCOPA | Latin | Indo-European |
| ja | Japanese | H | MGSM, XL-WiC | Kana & Hanzi | Japonic |
| ko | Korean | H | XL-WiC | Hangul | Koreanic |
| nl | Dutch | H | XL-WiC | Latin | Indo-European |
| qu | Quechua | L | XCOPA | Latin | Quechumaran |
| ro | Romanian | L | XQuAD | Latin | Indo-European |
| ru | Russian | H | MGSM, XQuAD | Cyrillic | Indo-European |
| sw | Swahili | L | MGSM | Latin | Niger–Congo |
| ta | Tamil | L | XCOPA | Tamil | Dravidian |
| te | Telegu | L | MGSM | Telegu | Dravidian |
| th | Thai | L | MGSM, XCOPA, XQuAD | Thai | Kra–Dai |
| tr | Turkish | H | XCOPA, XQuAD | Latin | Turkic |
| vi | Vietnamese | L | XCOPA, XQuAD | Latin | Austroasiatic |
| zh | Chinese | H | MGSM, XCOPA, XL-WiC XQuAD | Hanzi | Sino-Tibetan |
@@ C.1.1 Identifying Top-Activated Neurons
Each neuron in every MLP layer of the model is assigned a dedicated counter, initialized to . During vanilla evaluation, we monitor the activation of every neuron in the forward pass. Since LLMs typically employ ReLU-like Agarap (2018) activation functions (e.g., SwiGLU Shazeer (2020) for Llama series), a positive activation value can be interpreted as the neuron being “activated”. If a neuron is “activated”, we increment the corresponding counter by , otherwise no action. To ensure balanced input across our three datasets, we curated a COMBINED dataset, see §˜A.2 for details. After processing the inputs of a single ICL mode, each neuron accumulates an “activated” count. The neurons with the highest counts are identified as specialized neurons attributed to this ICL mode.